# Adversarial Confusion Attack: Disrupting Multimodal Large Language Models

### v2

@@ 1 Introduction
Most existing adversarial work targets classification errors, unsafe content steering, or jailbreak manipulation [madry2017towards, moosavidezfooli2017universal, qi2023visual, bailey2023image]. We address a distinct failure mode: confusion. A confusion attack aims to destabilize the model’s decoding process and produce high-confidence hallucinations or incoherent text, thereby preventing the model from forming a reliable understanding of the scene. We study a formulation of the confusion attack in which we maximize the next-token Shannon entropy of the model’s output distribution. This objective disrupts the decoder’s internal state and drives the model toward unstable token generation.
Prior work has shown that aligned multimodal models are vulnerable to universal attacks and patch-style perturbations [rahmatullaev2025universaladversarialattackaligned, aichberger2025attackingmultimodalosagents], that proprietary systems such as GPT-4 can be affected by adversarial examples [hu2025transferable, liu2025scalinglawsblackbox], and that perturbations and ensemble design follow scaling laws that govern black-box attack success [chen2024rethinking, liu2025scalinglawsblackbox, liu2017delving]. Our work complements recent adversarial research on MLLMs and contributes the following:
@@ 4 Discussion
Limitations & Future Work. This study uses an entropy-maximization objective implemented with PGD, a basic first-order adversarial optimization technique. Future work should investigate whether feature-level disruptions or more advanced momentum-based adversarial methods [chen2024rethinking] can help bridge the entropy gap between white-box and black-box settings. Enhancing robustness to compression, rendering, and small geometric transformations is also important for real-world deployment [athalye2018synthesizingrobustadversarialexamples]. The adversarial confusion attack also warrants evaluation within complex, multi-step agentic workflows [zhou2024webarena]. A particularly interesting direction is exploring how adversarial confusion can be embedded into website design, such as through background textures or UI color schemes.

### v3

@@ 1 Introduction
Most existing adversarial work targets classification errors, unsafe content steering, or jailbreak manipulation [madry2017towards, moosavidezfooli2017universal, qi2023visual, bailey2023image, ding2025practical, akiri2025safety, wang2025webinject, zhang2025realistic]. We address a distinct failure mode: confusion. A confusion attack aims to destabilize the model’s decoding process and produce high-confidence hallucinations or incoherent text, thereby preventing the model from forming a reliable understanding of the scene. We study a formulation of the confusion attack in which we maximize the next-token Shannon entropy of the model’s output distribution. This objective disrupts the decoder’s internal state and drives the model toward unstable token generation.
Prior work has shown that aligned multimodal models are vulnerable to universal attacks and patch-style perturbations [rahmatullaev2025universaladversarialattackaligned, aichberger2025attackingmultimodalosagents, hu2025c2, balakrishnan2025visor], that proprietary systems such as GPT-4 can be affected by adversarial examples [hu2025transferable, liu2025scalinglawsblackbox], and that perturbations and ensemble design follow scaling laws that govern black-box attack success [chen2024rethinking, liu2025scalinglawsblackbox, liu2017delving]. Our work complements recent adversarial research on MLLMs and contributes the following:
@@ 4 Discussion
Limitations & Future Work. This study uses an entropy-maximization objective implemented with PGD, a basic first-order adversarial optimization technique. Future work should investigate whether feature-level disruptions or more advanced momentum-based adversarial methods [chen2024rethinking, hu2025c2, balakrishnan2025visor] can help bridge the entropy gap between white-box and black-box settings. Enhancing robustness to compression, rendering, and small geometric transformations is also important for real-world deployment [athalye2018synthesizingrobustadversarialexamples]. The adversarial confusion attack also warrants evaluation within complex, multi-step agentic workflows [zhou2024webarena, ding2025practical, zhang2025realistic, wang2025webinject, wang2025advedm]. A particularly interesting direction is exploring how adversarial confusion can be embedded into website design, such as through background textures or UI color schemes.